Two identical metal spheres posses +60 c and -20 c of charges. They are bought in contact and then seperated by 10 cm.The force between them is
 

Grade:12

1 Answers

Khimraj
5 years ago
charge on each sphere after they are bought in contact = (60-20)/2 = 20 C
F = K*20*20/0.12 = 9*109*4*102/10-2 = 3.6*1014 N
Hope it clears.